OpenStudy (sydneeod):

Determine whether the sequence converges or diverges. If it converges, give the limit. 108, -54, 27, -27/2, ...

satellite73 (satellite73):

what do you multiply one number by to get the next?

OpenStudy (sydneeod):

-.5

satellite73 (satellite73):

ok good

OpenStudy (sydneeod):

so if r is between one and zero it is continuous?

satellite73 (satellite73):

so since \(|-5|<1\) the whole thing will go to zero hold on a sec

satellite73 (satellite73):

are you asking if the sequence converges, or if the sum converges? you have no plus signs there, just commas

satellite73 (satellite73):

Determine whether the \(\color{red}{ sequence}\) converges or diverges yes it does converge, the terms get smaller and smaller

OpenStudy (sydneeod):

i believe it's asking what the sum would be if the sequence does converge

satellite73 (satellite73):

by which i mean they go to zero this is nothing "continuous" about this, they are discrete numbers and as for \(r\) you are not being asked (as far as i can tell) to compute a sum, like \[\sum ar^n=\frac{a}{1-r}\] just the limit of the sequence

OpenStudy (sydneeod):

ooh okay okay

satellite73 (satellite73):

if they really want you to add, they should say find \[\sum108\times (-\frac{1}{2})^r\]

OpenStudy (sydneeod):

these are my answer choices, converges, 210 diverges converges, 540 converges, 0

OpenStudy (sydneeod):

we already know now that it doesn't converge

OpenStudy (sydneeod):

but i don't know how to find the limit of a sequence? I'm lost

OpenStudy (sydneeod):

@satellite73

OpenStudy (karim728):

this is a geometric series. with r =(-1/2) Note: it converges if |r|<1 and diverges if |r|>1. since r= |-1/2| =1/2 <1 it converges, now to find to what number it converges to, use this formula